# Finite basis Kolmogorov-Arnold networks: domain decomposition for data-driven and physics-informed problems

### v1

@@ Abstract
Kolmogorov-Arnold networks (KANs) have attracted attention recently as an alternative to multilayer perceptrons (MLPs) for scientific machine learning. However, KANs can be expensive to train, even for relatively small networks. Inspired by finite basis physics-informed neural networks (FBPINNs), in this work, we develop a domain decomposition method for KANs that allows for several small KANs to be trained in parallel to give accurate solutions for multiscale problems. We show that finite basis KANs (FBKANs) can provide accurate results with noisy data and for physics-informed training.
Keywords Kolmogorov-Arnold networks  $\cdot$ Physics-informed neural networks  $\cdot$ Domain decomposition
Figure 1: Graphical abstract.
@@ 1 Introduction
Scientific machine learning, predominantly developed using the theory of multilayer perceptrons (MLPs), has recently garnered a great deal of attention [1, 2, 3]. Kolmogorov-Arnold networks (KANs) [4] use the Kolmogorov-Arnold Theorem as inspiration for an alternative to MLPs. KANs offer advantages over MLPs in some cases, such as for continual learning and noisy data. While MLPs use trainable weights and biases on edges with fixed activation functions, KANs use trainable activation functions represented by splines. This switch may be beneficial because splines are easy to adjust locally and can switch between different resolutions. KANs have also been shown to provide increased interpretability over an MLP. Consequently, it is essential to provide a well-rounded analysis of the performance and applicability of KANs. An active area of research involves exploring how to utilize a KAN effectively and discerning in which cases a KAN might be preferred over an MLP. In this paper, we develop an architecture for KANs based on overlapping domain decompositions, which allows for more accurate solutions for complex problems.
@@
Since the publication of [4], numerous variations of KANs have been released, including fractional KANs [5], deep operator KANs [6], physics-informed KANs (PI-KANs)[7], KAN-informed neural networks (KINNs)[8], temporal KANs [9], graph KANs [10, 11, 12], Chebyshev KANs (cKANs) [13], convolutional KANs [14], ReLU-KANs [15], among others. KANs have been applied to satellite image classification [16], abnormality detection [17], and computer vision [18]. While these results are promising, there remains much to learn about how KANs can best be utilized and optimized. Many of the recent variations employ alternative functions to the B-spline parametrization, while others extend KANs through well-developed methods from the field of machine learning, such as physics-informed neural networks (PINNs) [19], deep operator networks (DeepONets) [20], and recurrent neural networks. Because KANs still follow the general network architecture of an MLP, save for the switch between the edges and neurons, it is often straightforward to apply such techniques to KANs.
@@
PINNs [19] have demonstrated success for a wide range of problems, including fluid dynamics [21, 22, 23, 24, 25], energy systems [26, 27, 28, 29, 30], and heat transfer [21]. PINNs allow for efficient and accurate solving of physical problems where robust data is lacking. This is accomplished by incorporating the governing differential equations of a system into the loss function of a neural network, which ensures that the resulting solution will satisfy the physical laws of the system. Although PINNs can successfully train with minimal data in many cases, there are notable cases in which challenges arise. For instance, PINNs can get stuck at the fixed points of a dynamical system while training [31]. PINNs are also difficult to train for multiscale problems, due to so-called spectral bias [32], or when the problem domain is large. Domain decomposition techniques can improve training in such cases [33, 34, 35, 36, 37]; for a general overview of domain decomposition approaches in scientific machine learning, see also [38, 39]. Rather than training the network over the entire domain, there are many approaches for breaking up the domain into subdomains and training networks on each subdomain. A popular choice for domain decomposition applied to PINNs is extended PINNs (XPINNs) introduced in [33], with a parallel version presented in [40], and extended in [34, 41]. XPINNs have different subnetworks for each domain, which are stitched together across boundaries and are trained in parallel. Another class of domain decomposition techniques for PINNs consists of time marching methods, where, in general, a network is initially trained for a small subdomain, and then the subdomain is slowly expanded until it contains the entire domain. Methods in this class include backward-compatible PINNs, which consist of a single network shared by all subdomains [35]. Also similar is the time marching approach in [36]. Finally, one particularly successful technique is finite basis PINNs (FBPINNs) [37, 42, 43, 44], which use partition of unity functions to weight the output of neural networks in each domain. In comparison to other techniques, FBPINNs offer simplicity because no boundary conditions are needed between adjacent domains. This work explores whether a similar technique can be used to enhance the training of a KAN for challenging cases.
@@
In addition to domain decomposition, many methods have been developed to improve the training of PINNs, including adaptive weighting schemes [45, 32], residual-based attention [46], adaptive residual point selection [47, 48, 49, 50, 51, 52], causality techniques [53], multifidelity PINNs [54, 55], and hierarchical methods to learn more accurate solutions progressively [56, 57, 58, 59, 60, 61, 62]. A substantial advantage of the methods presented in this work is that these additional techniques could be applied in combination with FBKANs to improve the training further.
@@
In this paper, we first introduce finite basis KANs (FBKANs) in section 2. In sections 3 and 4 we highlight some of the features of FBKANs applied to data-driven problems and physics-informed problems, respectively. We show that FBKANs can increase accuracy over KANs. One important feature of FBKANs is that the finite basis architecture serves as a wrapper around a KAN architecture. While we have chosen to focus on KANs as described in [4], most available extensions of KANs could be considered instead to increase accuracy or robustness further.
2 Methods
2.1 KANs
@@
In [4], the authors proposed approximating a multivariate function $f(\mathbf{x})$ by a model of the form
@@
denoted as a Kolmogorov-Arnold network. Here, $n_{l}$ is the number of layers in the KAN, $\{m_{j}\}_{j=0}^{n_{l}}$ is the number of nodes per layer, and $\phi_{i,j,k}$ are the univariant activation functions. We denote the right-hand side of eq. 1 as $\mathcal{K}(x)$. The activation functions are polynomials of degree $k$ on a grid with $g$ grid points. They are represented by a weighted combination of a basis function $b(x)$ and a B-spline,
$\phi(x)=w_{b}b(x)+w_{s}\text{spline}(x)$ (2)
@@
Here, $B_{i}(x)$ is a polynomial of degree $k$. $c_{i},w_{b},$ and $w_{s}$ are trainable parameters.
KANs evaluate the B-splines on a precomputed grid. In one dimension, a domain $[a,b]$ a grid with $g_{1}$ intervals has grid points $\{t_{0}=a,t_{1},t_{2},\ldots,t_{g_{1}}=b\}$; cf. [4]. Grid extension [4] allows for fitting a new, fine-grained spline to a coarse-grained spline, increasing the expressivity of the KAN. The coarse splines are transferred to the fine splines following the procedure in [4].
In this work, we do not consider the method for enforcing sparsity as outlined in [4] and instead only consider the mean squared error in the loss function. Many variations of KANs have been proposed recently. However, in this work, we consider only the formulation as outlined in [4], although we note that the domain decomposition method presented here would work with many variants of KANs.
@@ 2.2 Physics-informed neural networks
A PINN is a neural network (NN) that is trained to approximate the solution of an initial-boundary value problem (IBVP) by optimizing loss terms accounting for initial conditions, boundary conditions, and a residual term using backpropagation to calculate derivatives. In particular, the residual term represents how well the output satisfies the governing differential equation [19]. More specifically, we aim at approximating the solution $f$ of a generic IBVP
$f_{t}+\mathcal{N}_{x}[f]$ $=0,$ $x\in\Omega,t\in[0,T],$ (3)
$f(x,t)$ $=g(x,t),$ $x\in\partial\Omega,t\in[0,T],$
$f(x,0)$ $=u(x),$ $x\in\Omega,$
@@
over the open domain $\Omega\in\mathbb{R}^{d}$ with boundary $\partial\Omega.$ Here, $x$ and $t$ represent the spatial and temporal coordinates, respectively, $\mathcal{N}_{x}$ is the differential operator with respect to $x$, and $g$ and $u$ are given functions representing the boundary and initial conditions, respectively. The PINN model is then trained to minimize the mean squared error for the initial conditions, boundary conditions, and residual (physics) sampled at $N_{ic},N_{bc},N_{r}$ data points. We denote the sampling points for the initial conditions, boundary conditions, and residual as $\{x_{ic}^{i},u(x_{ic}^{i})\}_{i=1}^{N_{ic}}$, $\{(x_{bc}^{i},t_{bc}^{i}),g(x_{bc}^{i},t_{bc}^{i})\}_{i=1}^{N_{bc}}$, and $\{(x_{r}^{i},t_{r}^{i})\}_{i=1}^{N_{r}}$, respectively. We then optimize the following weighted loss function with respect to the trainable parameters $\theta$ of the NN model $f_{\theta}$:
$\mathcal{L}(\theta)$ $=\lambda_{ic}\mathcal{L}_{ic}(\theta)+\lambda_{bc}\mathcal{L}_{bc}(\theta)+\lambda_{r}\mathcal{L}_{r}(\theta),$ (4)
$\mathcal{L}_{ic}(\theta)$ $=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(f_{\theta}(x_{ic}^{i},0)-u(x_{ic}^{i})\right)^{2},$ (5)
$\mathcal{L}_{bc}(\theta)$ $=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(f_{\theta}(x_{bc}^{i},t_{bc}^{i})-g(x_{bc}^{i},t_{bc}^{i})\right)^{2},$ (6)
$\mathcal{L}_{r}(\theta)$ $=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\frac{\partial}{\partial t}f_{\theta}(x_{r}^{i},t_{r}^{i})+\mathcal{N}_{x}[f_{\theta}(x_{r}^{i},t_{r}^{i})]\right)^{2},$ (7)
@@
where $\lambda_{ic},\lambda_{bc},\lambda_{r}$ are weights for the different loss terms. The choice of these weights can have a significant impact on the training process and model performance. While adaptive schemes are available for choosing the weights [45, 32, 63, 64], in this work, we set the weights manually.
2.3 FBPINNs
@@
In FBPINNs [37, 42, 43, 44], we decompose the space domain $\Omega$ or the space-time domain $\Omega\times[0,T]$ into overlapping subdomains. Each overlapping subdomain $\Omega_{j}$ is the interior of the support of a corresponding function $\omega_{j}$, and all functions $\omega_{j}$ form a partition of unity. In particular, in the case of $L$ overlapping subdomains, we have
$\Omega=\bigcup_{j=1}^{L}\Omega_{j},\quad{\rm supp}(\omega_{j})=\overline{\Omega_{j}},\text{ and}\quad\sum_{j=1}^{L}\omega_{j}\equiv 1\text{ in }\Omega.$
@@
where $l=\max(x)-\min(x)$.
There are multiple ways to define the partition of unity functions. Here, we construct them based on the expression
$\omega_{j}=\frac{\hat{\omega}_{j}}{\sum_{j=1}^{L}\hat{\omega}_{j}},$ (8)
@@
$\hat{\omega}_{j}(x)=\begin{cases}1&L=1,\\ \left[1+\cos\left(\pi(x-\mu_{j})/\sigma_{j}\right)\right]^{2}&L>1,\end{cases}$ (9)
with $\mu_{j}=l(j-1)/(L-1)$ and $\sigma_{j}=(\delta l/2)/(L-1)$ representing the center and half-width of each subdomain, respectively. An example of the one-dimensional partition of unity functions with $L=4$ is depicted in fig. 2.
Figure 2: Example the partition of unity functions on the domain $\Omega=[0,2]$ with $L=4$ subdomains.
In multiple dimensions, $\mathbf{x}\in\mathbb{R}^{d}$, we then obtain
$\hat{\omega}_{j}(x)=\begin{cases}1&L=1,\\ \Pi_{i=1}^{d}\left[1+\cos\left(\pi(x_{i}-\mu_{ij})/\sigma_{ij}\right)\right]^{2}&L>1,\end{cases}$ (10)
with $\mu_{ij}$ and $\sigma_{ij}$ representing the center and half-width of each subdomain $j$ in each direction $i$. An example of of a single partition of unity function with $d=2$ and $L=4$ is shown in fig. 3.
@@
$f_{\theta}(x)=\sum_{j=1}^{L}\omega_{j}(x)f_{j,\theta^{j}}(x),$
where the $f_{j,\theta^{j}}$ is the neural network with parameters $\theta^{j}$ that corresponds to the subdomain $\Omega_{j}$; it is localized to $\Omega_{j}$ by multiplication with the partition of unity function $\omega_{j}$, which is zero outside of $\Omega_{j}$. The FBPINN model is trained in the same way as the PINN model, i.e., using initial, boundary, and residual loss terms; cf. section 2.2. Note that, in the original FBPINNs paper [37], hard enforcement of initial and boundary conditions is employed, such that only the residual loss term remains. Moreover, in [43], the approach has been extended to multilevel domain decompositions, and in [44], this domain decomposition approach has been applied to long-range time-dependent problems.
2.4 FBKANs
In FBKANs, we replace the NN used in FBPINNs with a KAN The function approximation in eq. 1 then becomes
$f(\mathbf{x})$ $\approx\sum_{j=1}^{L}\omega_{j}(x)\left[\sum_{i_{n_{l}-1}=1}^{m_{n_{l}-1}}\varphi^{j}_{n_{l}-1,i_{n_{l}},i_{n_{l}-1}}\left(\sum_{i_{n_{l}-2}=1}^{m_{n_{l}-2}}\ldots\left(\sum_{i_{1}=1}^{m_{1}}\varphi^{j}_{1,i_{2},i_{1}}\left(\sum_{i_{0}=1}^{m_{0}}\varphi^{j}_{0,i_{1},i_{0}}(x_{i_{0}})\right)\right)\ldots\right)\right]$ (11)
$=\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x;\theta^{j})$
where $\mathcal{K}^{j}(x;\theta^{j})$ denotes the $j^{th}$ KAN with trainable parameters $\theta^{j}$ and $\omega_{j}$ denotes the corresponding finite basis partition of unity function introduced in section 2.3. FBKANs are trained to minimize the loss function

### v2

@@ Abstract
Kolmogorov-Arnold networks (KANs) have attracted attention recently as an alternative to multilayer perceptrons (MLPs) for scientific machine learning. However, KANs can be expensive to train, even for relatively small networks. Inspired by finite basis physics-informed neural networks (FBPINNs), in this work, we develop a domain decomposition-based architecture for KANs that allows for several small KANs to be trained in parallel to give accurate solutions for multiscale problems. We denote this new approach as finite basis KANs (FBKANs) and show that it yields accurate results for noisy data as well as physics-informed training.
Keywords Kolmogorov-Arnold networks $\cdot$ Physics-informed neural networks $\cdot$ Domain decomposition $\cdot$ Network architecture $\cdot$ Multiscale problems $\cdot$ Finite basis PINNs
Figure 1: Graphical abstract: The computational domain is decomposed into overlapping subdomains, and an individual KAN model is defined on each subdomain. Then, the global model output is obtained by scaling the outputs of the local KAN models with partition of unity functions and summing up the local contributions.
@@ 1 Introduction
Scientific machine learning, predominantly developed using the theory of multilayer perceptrons (MLPs), has garnered a great deal of attention [karniadakis2021physics, baker2019workshop, carter2023advanced] over recent years. An alternative model architecture, which is based on the Kolmogorov-Arnold representation theorem [kolmogorov1957representations], has recently been introduced in [liu2024kan]. KANs offer advantages over MLPs in some contexts, such as for continual learning and noisy data. While MLPs use trainable weights and biases on edges with fixed activation functions, KANs employ trainable activation functions represented by splines. This switch may be beneficial because splines are easy to adjust locally and adapt to different resolutions. KANs have also been shown to provide increased interpretability over an MLP. Consequently, it is essential to provide a thorough analysis of the performance and applicability of KANs. An active area of research involves exploring how to utilize a KAN effectively and discerning in which cases KANs might be preferred over classical MLPs. In this paper, we develop an architecture for KANs based on overlapping domain decompositions, which allows for more accurate solutions for complex, and particularly multiscale, problems.
@@
Since the publication of [liu2024kan], numerous variations of KANs have been released, including fractional KANs [aghaei2025fkan], deep operator KANs [abueidda2025deepokan], physics-informed KANs (PI-KANs)[shukla2024comprehensive], KAN-informed neural networks (KINNs)[wang2025kolmogorov], temporal KANs [genet2024tkan], graph KANs [kiamari2024gkan, decarlo2024kolmogorovarnoldgraphneuralnetworks, bresson2024kagnnskolmogorovarnoldnetworksmeet], Chebyshev KANs (cKANs) [ss2024chebyshev], convolutional KANs [Bodner2024], ReLU-KANs [qiu2024relu], among others. KANs have been applied to satellite image classification [cheon2024kolmogorov], abnormality detection [huang2024abnormality], and computer vision [azam2024suitability]. Although these results are promising, many open questions remain about how KANs can be best utilized and optimized. Many of the recent variations employ alternative functions to the B-spline parametrization, while others extend KANs through well-developed methods from the field of machine learning, such as physics-informed neural networks (PINNs) [raissi2019physics], deep operator networks (DeepONets) [lu2019deeponet], and recurrent neural networks. Because KANs still follow the general network architecture of an MLP, aside from the switch between the edges and neurons, it is often straightforward to apply such techniques to KANs.
@@
PINNs [raissi2019physics] have demonstrated success for a wide range of problems, including fluid dynamics [cai2021physics, jin2021nsfnets, raissi2020hidden, mahmoudabadbozchelou2022nn, almajid2022prediction], energy systems [chen2023physics, misyris2020physics, huang2022applications, moya2023dae, bento2023physics], and heat transfer [cai2021physics]. PINNs allow for the efficient and accurate solution of physical problems where robust data is lacking. This is accomplished by incorporating the governing differential equations of a system into the loss function of a neural network, which ensures that the resulting solution will satisfy the physical laws of the system. Although PINNs can successfully train with little or no data in many cases, there are notable cases in which challenges arise. For instance, PINNs can get stuck at the fixed points of dynamical systems during training [rohrhofer2023ds]. PINNs are also difficult to train for multiscale problems, due to the so-called spectral bias [rahaman_spectral_2019, wang2022and], or when the problem domain is large. Domain decomposition techniques can improve training in such cases [jagtap2020extended, penwarden2023unified, mattey2022novel, wight2020solving, moseley2023finite]; for a general overview of domain decomposition approaches in scientific machine learning, see also [heinlein_combining_2021, klawonn_machine_2024]. Rather than training the network over the entire domain, there are many approaches for breaking up the domain into subdomains and training networks on each subdomain. A popular choice for domain decomposition applied to PINNs is extended PINNs (XPINNs) introduced in [jagtap2020extended], with a parallel version presented in [shukla2021parallel], and extended in [penwarden2023unified, hu2022extended]. XPINNs employ a separate subnetwork for each subdomain, which are stitched together across boundaries and trained in parallel. Another class of domain decomposition techniques for PINNs consists of time marching methods, where, in general, a network is initially trained for a small subdomain, and then the subdomain is slowly expanded until it contains the entire domain. Methods in this class include backward-compatible PINNs, which consist of a single network shared by all subdomains [mattey2022novel]. Also similar is the time marching approach in [wight2020solving]. Finally, a particularly successful technique is finite basis PINNs (FBPINNs) [moseley2023finite, dolean_finite_2024, dolean_multilevel_2024, heinlein2024multifidelity], which use partition of unity functions to weight the output of neural networks in each domain; these have also been extended to randomized neural networks [shang2025overlapping, anderson_elm-fbpinn_2024]. In comparison to other techniques, FBPINNs offer simplicity because the boundary conditions between adjacent domains do not have to be enforced via additional penalty terms in the loss function. This work explores whether a similar technique can be used to enhance the training of a KAN for challenging cases.
@@
In addition to domain decomposition, many methods have been developed to improve the training of PINNs, including adaptive weighting schemes [mcclenny2020self, wang2022and], residual-based attention [anagnostopoulos2024residual], adaptive residual point selection [wu2023comprehensive, mao2023physics, hou2023enhancing, lau2024pinnacle, nabian2021efficient, gao2023failure, visser_pacmann_2024], causality techniques [wang2022causality], multifidelity PINNs [penwarden2022multifidelity, meng2019multifidelity], and hierarchical methods to learn more accurate solutions progressively [howard2023continual, howard2023stacked, wang2024multi, ainsworth2021galerkin, ainsworth2022galerkin, trask2022hierarchical, aldirany2024multi]. A substantial advantage of the methodology presented in this work is that it can easily be combined with the aforementioned techniques to improve the training further.
@@
This paper is structured as follows: We first introduce finite basis KANs (FBKANs) in section˜2. In sections˜3 and 4, we highlight some of the features of FBKANs applied to data-driven and physics-informed problems, respectively. We show that FBKANs can increase accuracy over KANs. One important feature of FBKANs is that the finite basis architecture serves as a wrapper around a KAN architecture. While we have chosen to focus on KANs as described in [liu2024kan], most available extensions of KANs could be considered instead to increase accuracy or robustness further.
2 Methodology
2.1 Kolmogorov-Arnold Networks
@@
In [liu2024kan], the authors proposed approximating a multivariate function $f(\mathbf{x})$ by a model of the form
@@
which they denote as a Kolmogorov-Arnold network. Here, $n_{l}$ is the number of layers in the KAN, $\{m_{j}\}_{j=0}^{n_{l}}$ is the number of nodes per layer, $\varphi_{i,j,k}$ are the univariate activation functions, and $\mathbf{x}=(x_{1},\ldots,x_{m_{0}})$. We denote the right-hand side of eq.˜1 as $\mathcal{K}(\mathbf{x})$. The activation functions $\varphi_{k-1,i_{k},i_{k-1}}$ in the $k$th layer are polynomials of degree $k$ on a grid with $g$ grid points. They are represented by a weighted combination of a basis function $b(x)$ and a B-spline,
$\varphi(x)=w_{b}b(x)+w_{s}\text{spline}(x),$ (2)
@@
Here, $B_{i}(x)$ is a polynomial of degree $k$, and $c_{i}$, $w_{b}$, and $w_{s}$ are trainable parameters.
KANs evaluate the B-splines on a precomputed grid. In one dimension, in a domain $[a,b]$, a grid with $g_{1}$ intervals has grid points $\{t_{0}=a,t_{1},t_{2},\ldots,t_{g_{1}}=b\}$; cf. [liu2024kan]. Grid extension [liu2024kan] allows for fitting a new, fine-grained spline to a coarse-grained spline, increasing the expressivity of the KAN. The coarse splines are transferred to the fine splines following the procedure described in [liu2024kan] and [rigas2024adaptive].
In this work, we do not consider the method for enforcing sparsity of the trainable activation functions as outlined in [liu2024kan] and instead only consider the mean squared error (MSE) in the loss function. Many variations of KANs have been proposed recently. However, in this work, we consider only the formulation as outlined in [liu2024kan], although we note that the domain decomposition method presented here could be combined with many variants of KANs.
@@ 2.2 Physics-informed neural networks
A physics-informed neural network (PINN) is a neural network (NN) that is trained to approximate the solution of an initial-boundary value problem (IBVP) by optimizing loss terms accounting for initial conditions, boundary conditions, and a residual term using backpropagation to calculate derivatives. In particular, the residual term represents how well the output satisfies the governing differential equation [raissi2019physics]; for early approaches of combining differential equations with neural networks, we refer to [dissanayake_neural-network-based_1994, lagaris_artificial_1998]. More specifically, we aim at approximating the solution $f$ of a generic IBVP
$f_{t}+\mathcal{N}_{\mathbf{x}}[f]$ $=0,$ $x\in\Omega,t\in[0,T],$ (3)
$f(\mathbf{x},t)$ $=g(\mathbf{x},t),$ $x\in\partial\Omega,t\in[0,T],$
$f(\mathbf{x},0)$ $=u(\mathbf{x}),$ $x\in\Omega,$
@@
over the open domain $\Omega\in\mathbb{R}^{d}$ with boundary $\partial\Omega.$ Here, $x$ and $t$ represent the spatial and temporal coordinates, respectively, $\mathcal{N}_{x}$ is the differential operator with respect to $x$, and $g$ and $u$ are given functions representing the boundary and initial conditions, respectively.
The PINN model is then trained to minimize the MSE for the initial conditions, boundary conditions, and residual (physics) sampled at $N_{ic},N_{bc},N_{r}$ data points. We denote the sampling points for the initial conditions, boundary conditions, and residual as $\{x_{ic}^{i},u(x_{ic}^{i})\}_{i=1}^{N_{ic}}$, $\{(x_{bc}^{i},t_{bc}^{i}),g(x_{bc}^{i},t_{bc}^{i})\}_{i=1}^{N_{bc}}$, and $\{(x_{r}^{i},t_{r}^{i})\}_{i=1}^{N_{r}}$, respectively. Then, we optimize the following weighted loss function with respect to the trainable parameters $\theta$ of the NN model $f_{\theta}$:
$\mathcal{L}(\theta)$ $=\lambda_{ic}\mathcal{L}_{ic}(\theta)+\lambda_{bc}\mathcal{L}_{bc}(\theta)+\lambda_{r}\mathcal{L}_{r}(\theta),$
$\mathcal{L}_{ic}(\theta)$ $=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(f_{\theta}(\mathbf{x}_{ic}^{i},0)-u(\mathbf{x}_{ic}^{i})\right)^{2},$
$\mathcal{L}_{bc}(\theta)$ $=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(f_{\theta}(\mathbf{x}_{bc}^{i},t_{bc}^{i})-g(\mathbf{x}_{bc}^{i},t_{bc}^{i})\right)^{2},$
$\mathcal{L}_{r}(\theta)$ $=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\frac{\partial}{\partial t}f_{\theta}(\mathbf{x}_{r}^{i},t_{r}^{i})+\mathcal{N}_{\mathbf{x}}[f_{\theta}(\mathbf{x}_{r}^{i},t_{r}^{i})]\right)^{2},$
@@
where $\lambda_{ic},\lambda_{bc},\lambda_{r}$ are weights for the different loss terms. The choice of these weights can have a significant impact on the training process and model performance. While adaptive schemes are available for choosing the weights [mcclenny2020self, wang2022and, wang2022improved, howard2024conjugate], in this work, we set the weights manually to approximately scale the terms in the loss function to have equal magnitude. The values of the weights used are given in Appendix A.
2.3 Finite basis physics-informed neural networks
@@
In finite basis physics-informed neural networks (FBPINNs) [moseley2023finite, dolean_finite_2024, dolean_multilevel_2024, heinlein2024multifidelity], we decompose the spatial domain $\Omega$ or the space-time domain $\Omega\times[0,T]$ into overlapping subdomains. Each overlapping subdomain $\Omega_{j}$ is the interior of the support of a corresponding function $\omega_{j}$, and all functions $\omega_{j}$ form a partition of unity. In particular, in the case of $L$ overlapping subdomains, we have
[…]
@@
where $l=\max(x)-\min(x)$. Then, there are multiple ways to define the partition of unity functions. Here, we construct them based on the expression
$\omega_{j}=\frac{\hat{\omega}_{j}}{\sum_{j=1}^{L}\hat{\omega}_{j}},$
@@
$\hat{\omega}_{j}(x)=\begin{cases}1&L=1,\\ \left[1+\cos\left(\pi(x-\mu_{j})/\sigma_{j}\right)\right]^{2}&L>1,\end{cases}$
with $\mu_{j}=l(j-1)/(L-1)$ and $\sigma_{j}=(\delta l/2)/(L-1)$ representing the center and half-width of each subdomain, respectively. An example of the one-dimensional partition of unity functions with $L=4$ is depicted in fig.˜2.
Figure 2: Example of the partition of unity functions on the domain $\Omega=[0,2]$ with $L=4$ subdomains.
In multiple dimensions, $\mathbf{x}\in\mathbb{R}^{d}$, we then obtain the partition of unity functions as the tensor product of the one-dimensional functions:
$\hat{\omega}_{j}(\mathbf{x})=\begin{cases}1&L=1,\\ \Pi_{i=1}^{d}\left[1+\cos\left(\pi(x_{i}-\mu_{ij})/\sigma_{ij}\right)\right]^{2}&L>1,\end{cases}$
with $\mu_{ij}$ and $\sigma_{ij}$ representing the center and half-width of each subdomain $j$ in each dimension $i$. An example of an overlapping domain decomposition and of a single partition of unity function with $d=2$ and $L=4$ are shown in fig.˜3.
@@
$f_{\theta}(\mathbf{x})=\sum_{j=1}^{L}\omega_{j}(x)f_{j}(\mathbf{x};\theta^{j}),$ (4)
where $f_{j}(\cdot;\theta^{j})$ is the neural network with parameters $\theta^{j}$ that corresponds to the subdomain $\Omega_{j}$; it is localized to $\Omega_{j}$ by multiplication with the partition of unity function $\omega_{j}$, which is zero outside of $\Omega_{j}$. The FBPINN model is trained in the same way as the PINN model, i.e., using initial, boundary, and residual loss terms; cf. section˜2.2. Note that, in the original FBPINNs paper [moseley2023finite], hard enforcement of initial and boundary conditions is employed, such that only the residual loss term remains. Moreover, in [dolean_multilevel_2024] the approach has been extended to multilevel domain decompositions, in [heinlein2024multifidelity] this domain decomposition approach has been applied to long-range time-dependent problems, and in [shang2025overlapping, anderson_elm-fbpinn_2024] it has been extended to randomized neural networks.
2.4 Finite basis Kolmogorov–Arnold networks
In finite basis Kolmogorov–Arnold networks (FBKANs), we replace the NN used in FBPINNs with a KAN. The function approximation in eq.˜1 then becomes
$f(\mathbf{x})$ $\approx\sum_{j=1}^{L}\omega_{j}(x)\left[\sum_{i_{n_{l}-1}=1}^{m_{n_{l}-1}}\varphi^{j}_{n_{l}-1,i_{n_{l}},i_{n_{l}-1}}\left(\sum_{i_{n_{l}-2}=1}^{m_{n_{l}-2}}\ldots\left(\sum_{i_{1}=1}^{m_{1}}\varphi^{j}_{1,i_{2},i_{1}}\left(\sum_{i_{0}=1}^{m_{0}}\varphi^{j}_{0,i_{1},i_{0}}(x_{i_{0}})\right)\right)\ldots\right)\right]$
$=\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}_{j}(\mathbf{x};\theta^{j}),$ (5)
where $\mathcal{K}_{j}(\mathbf{x};\theta^{j})$ denotes the $j^{th}$ KAN with trainable parameters $\theta^{j}$ and $\omega_{j}$ denotes the corresponding finite basis partition of unity function introduced in section˜2.3. FBKANs are trained to minimize the PINN loss function